OpenStudy (anonymous):

what is the sum of these mixed numbers 2 2/3 + 3 2/3 options are 6 1/3 ,5 1/3 ,7 1/3, 6 /3 /4

OpenStudy (anonymous):

6 1/3

OpenStudy (anonymous):

8/3+11/3=19/3=6 1/3